Please provide a step by step solution of how to complete this problem

Windhoek Mines, Ltd., of Namibia, is contemplating the purchase of equipment to exploit a mineral deposit on land to which the company has mineral rights. An engineering and cost analysis has been made, and it is expected that the following cash flows would be associated with opening and operating a mine in the area:

Cost of new equipment and timbers $ 400,000
Working capital required $ 220,000
Annual net cash receipts $ 155,000 *
Cost to construct new roads in year three $ 64,000
Salvage value of equipment in four years $ 89,000

*Receipts from sales of ore, less out-of-pocket costs for salaries, utilities, insurance, and so forth.

The mineral deposit would be exhausted after four years of mining. At that point, the working capital would be released for reinvestment elsewhere. The company’s required rate of return is 18%.

Click here to view Exhibit 13B-1 and Exhibit 13B-2, to determine the appropriate discount factor(s) using tables.

Required:

a. What is the net present value of the proposed mining project?

b. Should the project be accepted?
Answer #1

a. Net Present value Calculation

Particulars Amount Discount years PV factor @ 18% Detail calculation Net present value
Initial Cost(Cost of equipment) -4,00,000 0        1.00 (400000)*1/(1+18%)^0 = -4,00,000
Annual Net Cash flow receipts year-1 1,55,000 1        0.85 155000*1/(1+18%)^1 = 1,31,356
Annual Net Cash flow receipts year-2 1,55,000 2        0.72 155000*1/(1+18%)^2 = 1,11,319
Annual Net Cash flow receipts year-3 1,55,000 3        0.61 155000*1/(1+18%)^3 =      94,338
Annual Net Cash flow receipts year-4 1,55,000 4        0.52 155000*1/(1+18%)^4 =      79,947
Cost of construction of road in 3rd year     -64,000 3        0.61 (64000)*1/(1+18%)^3 =     -38,952
Salvage value (Terminal value)      89,000 4        0.52 89000*1/(1+18%)^4 =      45,905
Net present value      23,912

Changes in working capital are an integral component in calculating net cash flow (receipts from sale of ore, less out-of-pocket costs for salaries, utilities, insurance).

Revenue expenditure has already being considered while calculating net cash flow receipts from sale of ore.

b. Since NPV (Rs. 23,912) is positive, the project can be accepted at opportunity cost (18%)